An electron, relased in a region where the electric field is uniform, is observed to have an acceleration of 7.0e14 m/s^2 in the positive x direction.

(A) Determine the electric field producing the acceleration.

I found this to be 3.99e3 N/C that it points towards -x.

(B) Assuming the electron is released from rest, determine the time required for it to reach a speed of 11,200 m/s, the escape speed from Earth's surface.
Answer #1

given,

acceleration = 7.0 * 10^14 m/s^2

mass of electron = 9.11 * 10^-31 kg

force = mass * acceleration

also,

force = charge * electric field

so,

mass * acceleration = charge * electric field

9.11 * 10^-31 * 7.0 * 10^14 = 1.6 * 10^-19 * electric field

electric field = 3985.625 N/C

initial speed = 0 m/s

final speed = 11200 m/s

by first equation of motion

v = u + at

11200 = 0 +  7.0 * 10^14 * time

time = 1.6 * 10^-11sec